OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is even or the sum of the dice is greater than 6? (A 6 6 table of dice outcomes will help you to answer this question.)

OpenStudy (k8lyn911):

21 out of 36 outcomes give you a number greater than 6, and 5 out of 36 give you 6. So \[\frac{ 21 +5 }{ 36 } = \frac{ 26 }{ 36 } = \frac{ 13 }{ 18 }\]